40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Mon Jul 2 18:27:25 2012, yuta, Update, General, BS and ITMX chambers closed 
    Reply  Tue Jul 3 17:23:50 2012, steve, Update, VAC, pumpdown completed pd72.pngpd72vacnormal.pngpd72at9hrs.png
       Reply  Fri Jul 6 09:27:40 2012, steve, Update, VAC, 4 days at atm vent_moves.png
          Reply  Fri Jul 6 09:52:35 2012, steve, Update, VAC, pump down hick up d3wFireworks.pngpdhickup.png
Message ID: 6903     Entry time: Mon Jul 2 18:27:25 2012     Reply to this: 6906
Author: yuta 
Type: Update 
Category: General 
Subject: BS and ITMX chambers closed 

[Koji, Steve, Jamie, Jenne, Yuta]

We opened BS and ITMX chambers, took lots of photos, and closed them with heavy doors.
I turned off high voltage power supplies for PZTs and blocked PSL beam. We are ready for the pumping tomorrow.

Important photos we took:
  - positions of green optics at BS chamber, which was moved on the vent on Aug 2011
  - positions of PZT mirrors and cable connectors at BS chamber, which will be replaced with tip-tilts on the next vent
  - arrow on PR2 pointing HR (it was correct)
  - tried to take photos of clipping IR beam at BS OSEM holder from ITMX chamber
 
 We also took bunch of other photos.


Beam dump needed at BS chamber:
  We also checked some un-dumped beams at BS chamber. We need dumps;
  - behind MMT1, for unwanted transmitted beam
  - behind IPPOSSM3, for unwanted transmitted beam (IPPOSSM3 is the last mirror in BS chamber for IPPOS)

ELOG V3.1.3-